# Language Model Preference Evaluation with Multiple Weak Evaluators

### v1

@@ 6 Experiments on Instruct Tuning
In this section, we explore the effects of various data selection methods for model alignment on Llama-2-7B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023) through instruct tuning. Specifically, we randomly sampled 5000 data points from UltraFeedback (Cui et al., 2023), and utilized Qwen1.5-14B (Yang et al., 2024a) to generate 8 responses for each data point as instruct data. We then applied four different methods—Random, Longest (Zhao et al., 2024), ContraSolver (Zhang et al., 2024b), and our proposed GED—to select a subset of these responses for model alignment training. The Origin refers to the performance of the base model without alignment. The resulting models were evaluated on the HH-RLHF (Bai et al., 2022) benchmark, which consists of four sub-sets: Harmless (base), Helpful (base), Helpful (online), and Helpful (rejection). For evaluation, we adopted the same Reward model, following prior work (Song et al., 2024; Yu et al., 2023), to measure human preference levels gained by the models. These results are summarized in Table 3. To ensure a comprehensive evaluation, we also tested the models, using the Llama-2-7B backbone, across additional benchmarks such as LIMA (Zhou et al., 2023), Vicuna (Chiang et al., 2023), Koala (Vu et al., 2023), WizardLM (Xu et al., 2023) and Self-Instruct (Wang et al., 2022), in line with recent works (Chen et al., 2023; Zhang et al., 2024a; Hu et al., 2024).
@@ 7 Related Work
Reference-free evaluation metrics have a long history (Louis & Nenkova, 2013), which evaluates the generated text based on intrinsic properties and coherence with the context. Although they achieve high accuracy on matching inner-evaluator, the achievement suffers from spurious correlations such as perplexity and length (Durmus et al., 2022). Recently, people have started using a strong model (e.g., GPT-4) as an evaluator to perform a zero-shot reference-free evaluation on the weak models (Shen et al., 2023; Dubois et al., 2024b; Chen et al., 2023). However, using LLM-based preference evaluations can introduce inconsistencies in preference graphs, often resulting in cyclic preferences or contradictions when comparing multiple outputs.
@@
The concept of weak-to-strong supervision originates from the need to leverage noisy or partial labels in machine learning tasks, enabling the development of more robust models from imperfect data (Ratner et al., 2016; Zhang et al., 2023b; 2022). In LLMs, weak-to-strong supervision aids AI alignment by allowing weaker models to improve strong ones, enhancing performance without extensive data and supporting scalable oversight (Zheng et al., 2024; Guo & Yang, 2024; Tong et al., 2024). Similarly, in task-oriented LLMs, weak-to-strong learning improves LLM’s ability by enabling strong models to refine their data autonomously, boosting performance without extensive high-quality input (Zhang et al., 2023a; Yang et al., 2024b). Through weak-to-strong supervision, LLM performance can be significantly improved by iteratively transforming low-quality labels into more reliable ones, leading to more effective model training and robust outputs (Zakershahrak & Ghodratnama, 2024; Lang et al., 2024).
@@ A.1 Experimental setup
All experimental procedures were conducted on a machine equipped with an AMD EPYC 7543 32-Core Processor, 512GB memory, 128 CPUs, and four 80GB NVIDIA A800 GPUs. The code is available at https://github.com/ppsmk388/GED. The references to Llama-2-7B, Llama3-70B, Llama3-8B, Mistral-7B, Qwen2-7B, and Qwen2-72B in the main text refer to the specific models: Llama-2-7b-chat-hf, Meta-Llama-3-70B-Instruct, Meta-Llama-3-8B-Instruct, Mixtral-8x7B-Instruct-v0.1, Qwen2-7B-Instruct, and Qwen2-72B-Instruct. We utilized the reward model oasst-rm-2-pythia-6.9b-epoch-1 following prior works (Song et al., 2024; Yu et al., 2023). Each experiment was repeated three times, and the average performance was reported as the final result. The training was configured with a batch size of 1 per device, gradient accumulation steps of 4, a learning rate of 1e-5, and he model was trained for 3 epochs, with warmup over 20 steps and a cosine learning rate scheduler. For tasks such as AlpacaEval (Dubois et al., 2024b), we used GPT-4-o unless stated otherwise.
A.2 Dataset
@@
Appendix B Construction of the Preference Graph
@@
Appendix C Proof of Theorem 1
@@
Appendix D Denoising of Preference Graphs for GED

### v2

@@ 6 Experiments on Instruct Tuning
In this section, we explore the effects of various data selection methods for model alignment on Llama-2-7B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023) through instruct tuning. Specifically, we randomly sampled 5000 data points from UltraFeedback (Cui et al., 2023), and utilized Qwen1.5-14B (Yang et al., 2024a) to generate 8 responses for each data point as instruct data. We then applied four different methods—Random, Longest (Zhao et al., 2024), ContraSolver (Zhang et al., 2024b), and our proposed GED—to select a subset of these responses for model alignment training. The Origin refers to the performance of the base model without alignment. The resulting models were evaluated on the HH-RLHF (Bai et al., 2022) benchmark, which consists of four sub-sets: Harmless (base), Helpful (base), Helpful (online), and Helpful (rejection). For evaluation, we adopted the same Reward model, following prior work (Song et al., 2024; Yu et al., 2023), to measure human preference levels gained by the models. These results are summarized in Table 3. To ensure a comprehensive evaluation, we also tested the models, using the Llama-2-7B backbone, across additional benchmarks such as LIMA (Zhou et al., 2023), Vicuna (Chiang et al., 2023), Koala (Vu et al., 2023), WizardLM (Xu et al., 2023) and Self-Instruct (Wang et al., 2022), in line with recent works (Chen et al., 2023b; Zhang et al., 2024a; Hu et al., 2024).
@@ 7 Related Work
Reference-free evaluation metrics have a long history (Louis & Nenkova, 2013), which evaluates the generated text based on intrinsic properties and coherence with the context. Although they achieve high accuracy on matching inner-evaluator, the achievement suffers from spurious correlations such as perplexity and length (Durmus et al., 2022). Recently, people have started using a strong model (e.g., GPT-4) as an evaluator to perform a zero-shot reference-free evaluation on the weak models (Shen et al., 2023; Dubois et al., 2024b; Chen et al., 2023b). However, using LLM-based preference evaluations can introduce inconsistencies in preference graphs, often resulting in cyclic preferences or contradictions when comparing multiple outputs.
@@
The concept of weak-to-strong supervision originates from the need to leverage noisy or partial labels in machine learning tasks, enabling the development of more robust models from imperfect data (Ratner et al., 2016; Zhang et al., 2023b; 2022). In LLMs, weak-to-strong supervision aids AI alignment by allowing weaker models to improve strong ones, enhancing performance without extensive data and supporting scalable oversight (Zheng et al., 2024a; Guo & Yang, 2024; Tong et al., 2024). Similarly, in task-oriented LLMs, weak-to-strong learning improves LLM’s ability by enabling strong models to refine their data autonomously, boosting performance without extensive high-quality input (Zhang et al., 2023a; Yang et al., 2024b). Through weak-to-strong supervision, LLM performance can be significantly improved by iteratively transforming low-quality labels into more reliable ones, leading to more effective model training and robust outputs (Zakershahrak & Ghodratnama, 2024; Lang et al., 2024).
@@ 8 Conclusion
Acknowledgement
This project was partially supported by an OpenAI “Superalignment Fast” grant.
@@ A.1 Experimental Setup
All experimental procedures were conducted on a machine equipped with an AMD EPYC 7543 32-Core Processor, 512GB memory, 128 CPUs, and four 80GB NVIDIA A800 GPUs. The references to Llama-2-7B, Llama3-70B, Llama3-8B, Mistral-7B, Mistral-8×7B, Qwen2-7B, and Qwen2-72B in the main text refer to the specific models: Llama-2-7b-chat-hf, Meta-Llama-3-70B-Instruct, Meta-Llama-3-8B-Instruct, Mixtral-7B-Instruct-v0.3, Mixtral-8x7B-Instruct-v0.1, Qwen2-7B-Instruct, and Qwen2-72B-Instruct. We utilized the reward model oasst-rm-2-pythia-6.9b-epoch-1 following prior works (Song et al., 2024; Yu et al., 2023). Each experiment was repeated three times, and the average performance was reported as the final result. Our training script was adapted from the example provided in LlamaFactory (Zheng et al., 2024b), specifically https://github.com/hiyouga/LLaMA-Factory/blob/main/examples/train_full/llama3_full_sft_ds3.yaml. The training was configured with a batch size of 1 per device, gradient accumulation steps of 4, a learning rate of 1e-5, and he model was trained for 3 epochs, with warmup over 20 steps and a cosine learning rate scheduler. For generating diverse responses from LLMs, we followed the configuration in Yuan et al. (2024), setting $T=0.7$ and $p=0.9$. For tasks such as AlpacaEval (Dubois et al., 2024b), we used GPT-4-o unless stated otherwise.
A.2 Details of Evaluator Selection Across Different Tasks
In this subsection, we provide more detailed information about the selection of evaluators across different tasks.
Response Selection.
In the response selection task, we evaluated both single models and ensembles of evaluators using the GED method. For single model evaluation, we assessed the standalone performance of Llama3-8B, Mistral-7B, Qwen2-7B, and Qwen2-72B on benchmarks such as HumanEval, AlpacaEval, MATH, GSM8k, and GAIA. This served two purposes: first, to establish the baseline performance of smaller models (Llama3-8B, Mistral-7B, Qwen2-7B) that are later combined using GED, and second, to compare against Qwen2-72B, where we tested a baseline approach by randomly selecting one response from the 10 it generated for each question. For single evaluator evaluation, each model (Llama3-8B, Mistral-7B, Qwen2-7B, and Qwen2-72B) was also used as an evaluator to rank responses generated by Qwen2-72B. Notably, Qwen2-72B acted as a self-evaluator, selecting the best response from its own generated outputs. Finally, for multiple evaluator evaluation, our GED method combined the evaluations of Llama3-8B, Mistral-7B, and Qwen2-7B to rank responses.
Model Ranking.
For the model ranking task, we selected larger models as evaluators to ensure alignment with rankings produced by GPT-4. Specifically, we used Llama3-70B, Qwen2-72B, Qwen1.5-72B, and Mistral-8×7B as single evaluators. This choice was guided by two factors: performance considerations and practical feasibility. Larger models generally produce more reliable rankings, closely aligning with GPT-4, and the AlpacaEval benchmark, containing 805 tasks, makes the computational cost of using larger models acceptable. In the multiple evaluator setting, our GED method aggregated the evaluations from these four larger models to produce robust and consistent rankings. The combination of these high-capacity models ensures that our approach yields rankings that are both accurate and consistent across tasks.
Instruction Tuning.
In the instruction tuning task, the objective was to perform data selection for model training. For this task, we employed Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators. These models were selected because they balance computational efficiency and evaluation performance, making them suitable for iterative instruction tuning processes. The evaluators’ pairwise preferences were aggregated using GED to identify the most appropriate responses for training. This approach ensured that the selected data pairs reflected a consensus among the evaluators while keeping computational costs manageable.
A.3 Definition of Weak Evaluators
In this work, ”weak evaluators” are defined by their tendency to produce noisy or inconsistent pairwise preferences, not by their overall model capacity. Even advanced models like GPT-4-o can generate preference graphs with cycles, indicating that preference inconsistency is a universal challenge, regardless of model scale. Evaluators such as Llama3-8B, Mistral-7B, Qwen2-7B, and even GPT-4-o are considered weak if their pairwise evaluations exhibit significant noise or conflicts.
A.4 Evaluation Settings: Single Model vs. Single Evaluator
The ”single model” setting evaluates each model’s (Llama3-8B, Mistral-7B, Qwen2-7B, Qwen2-72B) outputs directly on benchmarks like HumanEval, AlpacaEval, MATH, GSM8k, and GAIA, without selection or modification. In contrast, the ”single evaluator” setting uses these models to select the best response from ten candidates generated by Qwen2-72B, assessing their evaluation capability. The key difference is that the single model setting focuses on generation quality, while the single evaluator setting assesses evaluation ability.
A.5 Definition of Cycle Rate
The Cycle Rate is the percentage of preference graphs with at least one cycle, indicating inconsistency in pairwise comparisons. For example, if an evaluator produces cycles in 100 out of 164 graphs for the HumanEval dataset, the Cycle Rate is $\frac{100}{164}\times 100=60.97\%$. A lower Cycle Rate indicates greater consistency, while a higher rate suggests evaluator biases or difficulties with ambiguous comparisons. This metric helps assess the reliability of evaluators, such as GPT-4-o and GPT-4-o-mini, across different datasets.
A.6 Dataset
@@
A.7 Aggregation Process in GED Across Different Tasks
The GED implementation differs between response selection and model ranking tasks due to their objectives. In response selection, GED aggregates preference graphs from multiple evaluators for each question into a single DAG. From this DAG, a final ranking is derived, and the top-ranked answer is selected as the output. Notably, aggregation is performed only at the per-question level, without a rank aggregation across questions. In model ranking, GED involves two steps: first, aggregating the evaluators’ preference graphs into a DAG for each question to rank the models, and second, employing a rank ensemble method to aggregate these per-question rankings into a final overall model ranking across all questions. For details on the aggregation modeling, see Section 2.3, covering Response Selection, Model Ranking, and Model Alignment.
Appendix B Alternative Evaluator Configurations for Model Ranking
To address concerns about the computational cost and fairness of using 70B-level models as weak evaluators in the model ranking task, we conducted additional experiments using smaller, more comparable models. Table 4 summarizes the performance of these models, both individually and when combined using GED. The results show that GED outperforms individual evaluators even when using smaller 7B-scale models, achieving an average score of 62.70 compared to the best individual performance of 57.92 (Qwen2-7B with graph denoising). This demonstrates that the aggregation of smaller models through GED effectively enhances performance while reducing computational costs. These findings validate the versatility of GED, showing that it can provide robust and accurate rankings without relying solely on large-scale models.
Table 4: Performance comparison in the model ranking task using 7B-scale models as evaluators on the AlpacaEval dataset. Higher values indicate better performance. GED achieves robust performance even with smaller evaluators.
Model Weight Score Kemeny
Weighted
Kemeny
Pairwise
Majority
Weighted
Pairwise
Majority
Avg.
Llama3-8B 35.88 45.80 45.80 47.23 46.85 44.31
with graph denoising 37.44 47.54 47.54 48.92 48.18 45.92
Qwen2-7B 55.34 52.87 52.87 56.05 56.59 54.74
with graph denoising 56.05 57.43 57.43 59.32 59.41 57.92
Mistral-7B 49.90 53.74 53.74 58.06 57.87 54.66
Single evaluator with graph denoising 50.47 55.06 54.92 61.39 61.21 56.61
Multiple evaluator GED 57.59 61.14 61.14 67.17 66.46 62.70
Appendix C Impact of Evaluator Quantity on Denoising Quality
We investigated how the number of evaluators affects the denoising quality in GED by conducting experiments using different numbers of evaluators in the response selection setting. Specifically, we evaluated the performance of GED when aggregating preferences from two, three, and four evaluators. The evaluators used were Llama3-8B, Mistral-7B, Qwen2-7B, and Gemma-9B [^2] https://huggingface.co/google/gemma-2-9b-it.
Table 5: Performance comparison of GED with varying numbers of evaluators across five benchmarks. Increasing the number of evaluators enhances denoising quality, reflected in improved performance metrics.
Evaluators Set HumanEval AlpacaEval MATH GSM8k GAIA Avg.
Llama3-8B, Mistral-7B 69.21 31.87 74.97 86.92 12.51 55.10
Llama3-8B, Mistral-7B, Qwen2-7B 70.73 32.44 75.58 88.18 13.33 56.05
Llama3-8B, Mistral-7B, Qwen2-7B, Gemma-9B 70.98 32.87 75.91 88.75 13.46 56.39
As shown in Table 5, increasing the number of evaluators consistently improves the denoising quality of GED, as reflected by higher performance across all benchmarks. For instance, the average performance improves from 55.10% with two evaluators to 56.39% with four evaluators. This enhancement can be attributed to the diverse perspectives and complementary strengths of multiple evaluators, contributing to a more robust and accurate aggregation of preferences. These results highlight the importance of selecting a diverse and capable set of evaluators to enhance GED’s effectiveness. Incorporating more evaluators allows the denoising process to better identify and mitigate inconsistencies in the preference graphs, leading to improved overall performance in response selection tasks.
Appendix D Construction of the Preference Graph
@@
Appendix E Proof of Theorem 1
@@
Appendix F Impact of Evaluator Weighting on GED Performance
In our theoretical analysis, GED assumes equal weighting of edges in the preference graphs. However, in practical scenarios, evaluators may have varying levels of reliability or expertise. Incorporating evaluator-specific confidence scores or performance metrics could enhance the effectiveness of GED. To investigate this, we conducted experiments using a weighted version of GED, referred to as WeightGED, under the Response Selection setting. In WeightGED, we assigned weights to evaluators based on their individual performance on specific datasets. For example, on the GSM8k dataset, the response selection accuracies for Llama3-8B, Mistral-7B, and Qwen2-7B were 62.19%, 67.24%, and 61.58%, respectively. These accuracies were normalized to compute evaluator weights:
[…]
These weights were used to scale the contributions of each evaluator’s preferences during graph construction. We compared the performance of GED and WeightGED across multiple benchmarks, as presented in Table 6.
Table 6: Performance comparison of GED and WeightGED using 7B-scale evaluators across five benchmarks. WeightGED marginally outperforms GED, demonstrating the potential benefits of incorporating evaluator-specific weights.
Method HumanEval AlpacaEval MATH GSM8k GAIA Avg.
GED 70.73 32.44 75.58 88.18 13.33 56.05
7B Evaluators WeightGED 70.97 32.67 75.71 88.24 13.56 56.23
GED 73.21 59.87 82.49 86.43 16.27 63.65
GPT Evaluators WeightGED 74.52 61.71 83.93 87.84 17.32 65.06
As shown in Table 6, WeightGED achieves marginal but consistent improvements over GED across all benchmarks. For instance, using 7B-scale evaluators, WeightGED improves the average performance from 56.05% to 56.23%. Similarly, with GPT evaluators, the average performance increases from 63.65% to 65.06%. These results suggest that incorporating evaluator-specific weights based on performance metrics can enhance the effectiveness of GED. Furthermore, we conducted additional experiments using stronger evaluators such as GPT-3.5, GPT-4-o-mini, and GPT-4-o. The weights were computed based on their respective performance accuracies, following the same normalization procedure. The improvements observed with these evaluators reinforce the potential of weighting schemes in enhancing GED. In summary, our preliminary findings indicate that integrating evaluator-specific confidence scores or performance metrics is a promising direction for future work. Systematically designing and optimizing these weighting schemes could lead to more robust and accurate evaluation frameworks.
Appendix G Evaluating GED on Diverse Metrics
To address the need for more comprehensive assessments beyond accuracy-based metrics, we expanded our evaluation of GED to include nuanced quality aspects of LLM-generated outputs, such as factuality, relevance, coherence, informativeness, helpfulness, and validity. This evaluation was motivated by the understanding that tasks involving LLMs often require subtle judgments beyond simple accuracy, and GED’s adaptability to these scenarios is crucial. Building upon this foundation, we conducted experiments following (Chen et al., 2023a). Specifically, we used Llama3-70B to generate ten candidate responses for each query, and GED was applied in the response selection setting with Llama3-8B, Mistral-7B, and Qwen2-7B serving as evaluators. Each metric—factuality, relevance, coherence, informativeness, helpfulness, and validity—was assessed to measure GED’s ability to enhance the overall quality of LLM outputs. The results, shown in Table 7, highlight GED’s effectiveness. GED consistently outperformed individual evaluators and random selection across all metrics. For instance, GED improved factuality by approximately 5 percentage points over the best individual evaluator. Similarly, it enhanced relevance and coherence, indicating better alignment with the query and logically consistent responses. GED also demonstrated notable improvements in informativeness and helpfulness, suggesting that the responses selected by GED provide more valuable and user-centric information. These findings confirm GED’s adaptability to tasks requiring nuanced quality assessments. By aggregating preferences from multiple evaluators, GED captures subtle qualities of generated content, enhancing not only accuracy but also the overall reliability and utility of LLM outputs. This adaptability reinforces GED’s value in scenarios demanding comprehensive evaluations of language models.
Table 7: Performance comparison on nuanced quality metrics (in %). GED outperforms individual evaluators and random selection across Factuality, Relevance, Coherence, Inform., Helpful. and Validity metrics.
Method Factuality Relevance Coherence Inform. Helpful. Validity Avg.
Random 86.73 87.91 92.47 77.62 17.48 48.92 68.52
Llama3-8B 88.59 89.91 94.41 79.77 18.48 50.92 70.35
Mistral-7B 89.10 90.29 94.85 79.95 18.55 51.13 70.65
Qwen2-7B 89.25 90.44 95.03 80.09 18.58 51.21 70.77
GED 94.73 95.91 97.36 86.62 19.48 55.92 75.00
Appendix H Expanded Related Work on Preference Denoising for LLMs
This work situates itself within the broader field of denoising preference evaluations for LLMs, addressing inconsistencies and noise in preference graphs. We acknowledge that existing literature has explored two primary approaches to this challenge: within-model preference modeling and modular pre-processing. Below, we provide a detailed comparison of GED with representative methods from these approaches.
Within-model approaches, such as robust DPO (rDPO) (Chowdhury et al., 2024) and conservative DPO (cDPO) (Mitchell, 2023), focus on integrating denoising mechanisms directly within the preference modeling process. These methods incorporate regularization techniques to mitigate the effects of noisy or adversarial preference data during model alignment. While these approaches are effective in refining the preference modeling pipeline, they are tightly coupled with specific models and tasks, limiting their versatility. In contrast, GED is a modular framework that operates as a pre-processing step, denoising preference graphs before downstream tasks, making it adaptable to a wide range of applications and models. Modular approaches like CURATRON (Naresh et al., 2024) address noise and missing comparisons in preference datasets using techniques such as low-rank matrix completion. While CURATRON effectively mitigates certain types of noise, it does not explicitly target cyclic inconsistencies (e.g., A $\succ$ B, B $\succ$ C, C $\succ$ A), which are a critical focus of GED. By leveraging a graph-based framework, GED detects and removes such cycles through its denoising process, ensuring that the resulting preference graph is acyclic and thus more consistent and reliable for downstream use. Additionally, GED distinguishes itself by providing theoretical guarantees for recovering the ground truth preference DAG under certain conditions. Furthermore, the ensemble mechanism in GED demonstrates the novel insight that combining weaker evaluators can surpass the performance of a single stronger evaluator, a feature not emphasized in the aforementioned methods.
Appendix I Impact of Evaluator Selection on GED Performance
The selection of weak evaluators is a critical factor influencing the performance of GED. While our initial experiments demonstrated the benefits of combining weaker evaluators, further analysis is necessary to understand how their diversity and capabilities affect GED’s effectiveness. To address this, we conducted additional experiments evaluating three different configurations of evaluator sets: the original set comprising Llama3-8B, Mistral-7B, and Qwen2-7B; a diverse set that adds Gemma-9B [^4] https://huggingface.co/google/gemma-2-9b-it to introduce more variation in model architecture and training data; and a higher-capability set that replaces smaller models with larger ones, specifically Llama3-70B, Mistral-8×7B, and Qwen2-72B. The results, shown in Table 8, reveal two key insights. First, incorporating diversity by adding Gemma-9B improves GED’s performance slightly, increasing the average score from 56.05% to 56.39%. This suggests that models with diverse training data and architectures can contribute to more robust aggregated evaluations. Second, replacing smaller evaluators with higher-capability models yields a more substantial improvement, with the average score rising to 60.89%. Notably, benchmarks like AlpacaEval and GAIA benefit the most from the advanced reasoning and language understanding capabilities of larger models. These findings demonstrate the importance of both diversity and capability in evaluator selection. Diversity provides marginal gains by bringing varied perspectives, while higher-capacity models contribute to more significant improvements by enhancing the overall quality of evaluations. This suggests that, when computational resources allow, incorporating advanced models into the evaluator set can meaningfully boost GED’s performance.
Table 8: Performance comparison with different evaluator sets across five benchmarks. The results highlight the impact of evaluator diversity and capability on GED’s effectiveness.
Evaluators Set HumanEval AlpacaEval MATH GSM8k GAIA Avg
Original 70.73 32.44 75.58 88.18 13.33 56.05
Diverse 70.98 32.87 75.91 88.75 13.46 56.39
Higher Capability 74.73 47.92 75.58 90.04 16.21 60.89
Appendix J Cost Considerations and Active Learning with ActiveGED
Using multiple evaluators and aggregating their preferences can be computationally expensive, especially when constructing dense preference graphs. This limitation becomes more pronounced in scenarios where preferences across all pairs of evaluators are required, which scales quadratically with the number of responses or models being compared. To address this, we clarify the computational trade-offs and propose an active learning-based approach to reduce the number of required pairwise evaluations. To further reduce the number of pairwise evaluations needed, we developed an active learning algorithm called ActiveGED. This algorithm strategically selects the most informative pairs for evaluation, effectively lowering the overall computational cost while maintaining high performance. ActiveGED combines random sampling with uncertainty-based selection to maximize information gain from each additional pairwise evaluation. We evaluated ActiveGED under budget constraints of 30% and 50% of the total possible pairwise comparisons, where we set the random budget ratio as 0.5. The results, presented in Table 9, demonstrate that ActiveGED achieves competitive performance with significantly fewer evaluations. For example, under a 50% budget, ActiveGED retains most of the performance benefits of full GED while cutting the number of pairwise comparisons in half.
Table 9: Performance comparison of ActiveGED under different budget constraints. ActiveGED achieves competitive performance with significantly fewer pairwise evaluations.
Evaluators Set of GED HumanEval AlpacaEval MATH GSM8k GAIA Avg.
Random 57.93 29.58 72.75 84.67 11.52 51.29
ActiveGED (30%) 67.28 30.96 74.65 85.73 11.39 54.00
ActiveGED (50%) 68.62 31.91 74.87 87.06 12.08 54.91
GED (Full Budget) 70.73 32.44 75.58 88.18 13.33 56.05
The algorithm behind ActiveGED is outlined in Algorithm 3. It begins by randomly selecting a portion of the budget to initialize the preference graph and then iteratively selects the most informative edges based on uncertainty, as estimated using PageRank scores. This approach balances exploration (random sampling) and exploitation (uncertainty-based selection) to efficiently construct an accurate preference graph.
Algorithm 3 ActiveGED
1:Candidate set $V=\{v_{1},v_{2},\dots,v_{n}\}$; Initial preference graph $G=(V,A,w)$; Evaluator set $E=\{\text{ev}_{1},\dots,\text{ev}_{k}\}$; Total budget $B$; Random budget ratio $\alpha\in(0,1)$
2:Updated preference graph $G^{*}=(V,A^{*},w^{*})$
3:Randomly select $m_{\text{rand}}=\alpha B$ edges from $(V\times V)\setminus A$ to form $A_{\text{rand}}$
4:for each edge $(u,v)\in A_{\text{rand}}$ do
5:     for each evaluator $\text{ev}_{i}\in E$ do
6:         Obtain preference weight $w_{i}(u,v)$
7:     end for
8:     Aggregate weights $w(u,v)=\frac{1}{k}\sum_{i=1}^{k}w_{i}(u,v)$
9:end for
10:Update $A^{*}\leftarrow A\cup A_{\text{rand}}$
11:Set current budget $b\leftarrow m_{\text{rand}}$
12:Compute PageRank $PR(v)$ for all $v\in V$ using $G^{*}=(V,A^{*},w^{*})$
13:while $b<B$ do
14:     for each unevaluated edge $(u,v)\in(V\times V)\setminus A^{*}$ do
15:         Estimate uncertainty $U(u,v)$ based on current PageRank scores
16:     end for
17:     Select edge $(u^{*},v^{*})$ with highest $U(u,v)$
18:     for each evaluator $\text{ev}_{i}\in E$ do
19:         Obtain preference weight $w_{i}(u^{*},v^{*})$
20:     end for
21:     Aggregate weights $w(u^{*},v^{*})=\frac{1}{k}\sum_{i=1}^{k}w_{i}(u^{*},v^{*})$
22:     Update $A^{*}\leftarrow A^{*}\cup\{(u^{*},v^{*})\}$
23:     Increment $b\leftarrow b+1$
24:     Recompute PageRank $PR(v)$ for all $v\in V$ using the updated $G^{*}$
25:end while
26:return $G^{*}=(V,A^{*},w^{*})$
ActiveGED demonstrates that by carefully selecting the most informative pairs, it is possible to achieve competitive performance with significantly fewer evaluations. This makes GED more scalable and practical for real-world applications where computational resources are limited.
Appendix K Mitigating Evaluator Biases in GED
Evaluator biases, such as position bias and token bias, can significantly impact the accuracy and fairness of the preference graphs used in GED. Addressing these biases is crucial for ensuring reliable evaluations and robust performance. In this section, we describe the strategies employed in our framework to mitigate these biases and discuss potential areas for future improvement. Position bias arises when evaluators exhibit a preference for a particular position in a pairwise comparison, such as consistently favoring the first or second option regardless of content. To counter this, we explicitly include both orderings of each question and its candidate answers in the response selection setting. Specifically, for a question $Q$ with two candidate answers $A_{1}$ and $A_{2}$, we evaluate both configurations:
  • Order 1:
        Question: [Question Text]
        Answer 1: [Answer Text A1]
        Answer 2: [Answer Text A2]
        Which one is better?
  • Order 2:
        Question: [Question Text]
        Answer 1: [Answer Text A2]
        Answer 2: [Answer Text A1]
        Which one is better?
By testing both (Q, A1, A2) and (Q, A2, A1), we ensure that any positional preferences of the evaluators are balanced out when constructing the preference graph. This approach minimizes the impact of position bias on the overall rankings. Token bias occurs when an evaluator exhibits a latent preference for a specific option or label (e.g., consistently favoring ”Option A” over ”Option B”). Our framework implicitly mitigates token bias through the aforementioned position-swapping strategy, which prevents evaluators from associating a fixed label with a specific position. By averaging the results across both orderings, any systematic preference for a particular label is effectively neutralized.
Appendix L Denoising of Preference Graphs for GED